OpenStudy (ntgarrett33):

how many people are in the US if 1/4th of the people equal 57.7 million

sam (.sam.):

(1/4)x=57.7 x=4(57.7)=230.8million

OpenStudy (ntgarrett33):

Thank You! :-)

OpenStudy (anonymous):

If 1/4 = 57.7m Then, you would have to multiply 57.7*4 57,700,700*4 = 230,800,000